OpenStudy (kaitlyn_nicole):

Write the equation of the line that passes through (1, 5) and (–2, 14) in slope-intercept form.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ 1}}\quad ,&{\color{blue}{ 5}})\quad % (c,d) &({\color{red}{ -2}}\quad ,&{\color{blue}{ 14}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (jdoe0001):

find the slope use either point's coordinate values plug them in the point-slope intercept solve for "y" be happy eat ice-cream

OpenStudy (kaitlyn_nicole):

okay, so y-5=3(x-1) ??

OpenStudy (jdoe0001):

what did you get for the slope anyway?

OpenStudy (kaitlyn_nicole):

\[\frac{ 14-5 }{ -2-1 }\] so that simplifies to \[\frac{ 9 }{ -3 }\] so the slope is -3

OpenStudy (jdoe0001):

yeap that's correct, tis -3 :) so y -5 = \(\bf -3\) (x-1) then solving for "y" y = -3x+3+5 y = -3x+8

OpenStudy (kaitlyn_nicole):

o so thats the slope intercept form?

OpenStudy (jdoe0001):

y = -3x + 8 ^ ^ slope y-intercept slope = -3 y-intercept = +8 thus, is called the slope-intercept form :)

OpenStudy (kaitlyn_nicole):

so what about point-slope form?

OpenStudy (kaitlyn_nicole):

ill make another question so you can get metals ;p

OpenStudy (jdoe0001):

the point slope form is right above :) \(\bf y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{}\\ \qquad \uparrow\\ \textit{point-slope form}\)